# DiaBlo: Diagonal Blocks Are Sufficient For Finetuning

### v1

@@ Results
The experimental results on HumanEval demonstrate that DiaBlo achieves strong performance in code generation on Llama3-8B. With 1.51% trainable parameters, DiaBlo reaches a Pass@1 score of 41.8% and a Pass@10 score of 62.9%, outperforming LoRA and DoRA across all metrics and closely matching the best-performing method, LoRI. With a reduced parameter budget of 0.76%, DiaBlo still maintains high performance, achieving a Pass@10 score of 61.9%. Notably, if trained for one additional epoch, DiaBlo further improves to a Pass@1 score of 42.65%, Pass@5 of 57.11%, and Pass@10 of 64.07%, which are comparable to the best results achieved by LoRI. On Mistral-7B, DiaBlo with $N=128$ achieves the best Pass@5 and Pass@10 scores and the second-best Pass@1 score, while DiaBlo with $N=64$ achieves the best Pass@1 score and the second-best Pass@5 and Pass@10 scores, outperforming LoRA, DoRA, and LoRI. In terms of HEx-PHI, DiaBlo outperforms all other methods with scores of 97.6% on Llama3-8B and 98.8% on Mistral-7B. These findings highlight DiaBlo’s effectiveness in both code generation and safety alignment, while maintaining a high level of parameter efficiency.
Table 4: Performance on code generation and safety alignment tasks. Results of baseline methods are taken from LoRI Zhang2025LoRIRC .
Model Method $\mathbf{r}/\mathbf{N}$Params HumanEval HEx-PHI
@@
Llama3-8B LoRA $r=32$ 1.12% 34.7 46.4 50.8 91.6
@@
Mistral-7B LoRA $r=32$ 1.25% 33.8 42.4 45.3 91.9
@@ 4.4 Fine-tuning with Quantized Models
To further reduce the memory footprint of fine-tuning, QLoRA guo2024lqlora and its variants store the frozen pretrained weights in low-precision while training small low-rank adaptation modules in higher precision. In this section, we explore the compatibility of the proposed DiaBlo with quantized models and evaluate its performance when fine-tuning with quantized models.
@@
We use DiaBlo to fine-tune Llama2-7B model in 4-bit and 2-bit, which are quantized by MagR zhang2024magr . Following the setting in ApiQ liao2024apiq , we fine-tune the model on Math10K hu-etal-2023-llm and then evaluate on test sets of AQuA ling2017program , GSM8K cobbe2021training , MAWPS koncel2016mawps , and SVAMP patel2021nlp .
@@ Results
We evaluate DiaBlo on 4-bit and 2-bit quantized models for arithmetic reasoning tasks and compare its performance with existing fine-tuning methods for quantized models, QLoRA dettmers2023qlora , GPTQ-LoRA GPTQLoRA , LoftQ li2023loftq , and ApiQ liao2024apiq . In the 4-bit setting, DiaBlo achieves the highest average accuracy of 54.8%, outperforming all other methods while also using significantly fewer trainable parameters. In the more challenging 2-bit setting, DiaBlo again outperforms existing baselines with an average accuracy of 48.7%. These results demonstrate that DiaBlo remains robust and effective under low-precision constraints, offering a highly parameter-efficient and accurate approach for fine-tuning quantized models. Notably, DiaBlo directly fine-tunes quantized models without relying on any customized quantization or initialization strategy. In contrast, state-of-the-art methods such as LoftQ li2023loftq and ApiQ liao2024apiq introduce specialized mechanisms to calibrate quantized weights and carefully initialize LoRA components to improve stability and performance. The simplicity of DiaBlo highlights its practical advantages while achieving superior or comparable results.
Table 5: Performance of fine-tuning quantized Llama2-7B model on arithmetic reasoning tasks. Results of baseline methods are taken from ApiQ liao2024apiq .
Method $\mathbf{r/N}$Params GSM8K SVAMP mawps AQuA AVG
LoRA $r=64$ 112M 43.6 59.4 85.0 27.0 53.7
@@
DiaBlo $N=64$ 70M 32.1 51.5 87.0 24.0 48.7
@@ 4.5 Memory and Computation Efficiency of DiaBlo
The simple structure of DiaBlo allows it to make use of efficient batched operations, helping reduce training time and memory usage. In this subsection, we evaluate the memory consumption and training speed of DiaBlo and compare them with other PEFT training methods.
@@
Table 6: Comparison of fine-tuning runtime.
Method $\mathbf{r/N}$Params Time (min)
LoRA $r=32$ 1.12% 170
DoRA $r=32$ 1.12% 480
DiaBlo $N=64$ 1.51% 170
DiaBlo $N=128$ 0.76% 170
To further validate this, we benchmark the fine-tuning speed of DiaBlo on the commonsense reasoning 170k dataset and compare it with LoRA and DoRA. The experiments are conducted on a single A100 GPU with 80GB memory, using a batch size of 8 and a gradient accumulation step of 4. The times for one-epoch fine-tuning are presented in Table 6. DiaBlo achieves training speeds exactly the same as LoRA, while DoRA takes significantly longer. This demonstrates the practical efficiency of the proposed DiaBlo method.

### v2

@@ Results
The experimental results are summarized in Table 3. DiaBlo achieves strong performance in the code generation task HumanEval on Llama3-8B. With 1.51% trainable parameters, DiaBlo reaches a Pass@1 score of 43.2% and a Pass@10 score of 63.5%, outperforming all other methods. Its Pass@5 score is also very cloes to the best result obtained by LoRI. With a reduced parameter budget of 0.76%, DiaBlo still maintains high performance, achieving a Pass@10 score of 61.9%. On Mistral-7B, DiaBlo with $N=128$ achieves the best Pass@5 and Pass@10 scores and the second-best Pass@1 score, while DiaBlo with $N=64$ achieves the best Pass@1 score and the second-best Pass@5 and Pass@10 scores, outperforming LoRA, DoRA, and LoRI. In terms of HEx-PHI, DiaBlo outperforms all other methods with scores of 97.6% on Llama3-8B and 98.8% on Mistral-7B. These findings highlight DiaBlo’s effectiveness in both code generation and safety alignment, while maintaining a high level of parameter efficiency.
Table 3: Performance on code generation and safety alignment tasks. Results of baseline methods are taken from LoRI (Zhang et al., 2025a).
Model Method $\mathbf{r}/\mathbf{N}$ #Params HumanEval HEx-PHI
@@
Llama3-8B Zero-shot N/A 0% 28.5 39.8 45.2 78.7
LoRA $r=32$ 1.12% 34.7 46.4 50.8 91.6
@@
Mistral-7B Zero-shot N/A 0% 28.6 38.2 41.9 80.5
LoRA $r=32$ 1.25% 33.8 42.4 45.3 91.9
@@ 4.4 Fine-tuning with Quantized Models
To further reduce the memory footprint of fine-tuning, QLoRA (Guo et al., 2024) and its variants store the frozen pretrained weights in low-precision while training small low-rank adaptation modules in higher precision. In this section, we explore the compatibility of the proposed DiaBlo with quantized models and evaluate its performance.
@@
We evaluate DiaBlo on quantized models. We fine-tune the quantized LLaMA2-7B and LLaMA2-13B models, which are quantized by GPTQ (Frantar et al., 2022) and MagR (Zhang et al., 2024). All models are fine-tuned on Math10K (Hu et al., 2023) and then evaluated on the test sets of AQuA (Ling et al., 2017), GSM8K (Cobbe et al., 2021), MAWPS (Koncel-Kedziorski et al., 2016), and SVAMP (Patel et al., 2021).
@@ Results
We evaluate DiaBlo on 4-bit and 2-bit quantized Llama2-7B and Llama2-13B models. The results are presented in Table 4. For the Llama2-7B model, MagR-DiaBlo achieves the highest average accuracy of 54.8% in the 4-bit setting. In the more challenging 2-bit setting, it again outperforms existing baselines with an average accuracy of 48.7%. The results on the larger Llama2-13B model further demonstrate DiaBlo’s scalability and robustness. Its advantage is particularly stark in the ultra-low 2-bit setting, where it achieves an average score of 55.1%, substantially outperforming baselines. The LoftQ and ApiQ methods jointly determine quantized weights and LoRA initialization to enhance fine-tuning performance. In contrast, DiaBlo directly fine-tunes quantized models without any customized quantization procedure or special initialization, highlighting its practical simplicity while still achieving superior or comparable results. To isolate the impact of quantization methods, we also evaluate GPTQ-DiaBlo and MagR-LoRA under the 2-bit setting. Under both GPTQ quantization and MagR quantization, simply replacing LoRA with DiaBlo leads to substantial performance improvements. This indicates that the observed gains come primarily from the fine-tuning method DiaBlo rather than from changes in the quantization method.
Table 4: Performance of fine-tuning quantized Llama2-7B and Llama2-13B models on arithmetic reasoning tasks. Results of baseline methods are taken from ApiQ (Liao et al., 2024).
Llama2-7B
Method $\mathbf{r/N}$ #Params GSM8K SVAMP MAWPS AQuA AVG
@@
GPTQ-DiaBlo $N=64$ 70M 33.3 50.9 84.0 22.8 47.8
MagR-DiaBlo $N=64$ 70M 32.1 51.5 87.0 24.0 48.7
Llama2-13B
4-bit QLoRA $r=64$ 239M 54.8 69.4 87.0 26.8 59.5
GPTQ-LoRA $r=64$ 239M 53.2 67.5 85.3 25.6 57.9
LoftQ $r=64$ 239M 54.9 66.5 87.7 23.9 58.3
ApiQ-bw $r=64$ 239M 55.3 67.4 87.8 25.6 59.0
MagR-DiaBlo $N=64$ 189M 55.5 66.9 89.9 26.0 59.6
2-bit QLoRA $r=64$ 239M 0.5 0.7 0.1 0.9 0.6
GPTQ-LoRA $r=64$ 239M 31.9 49.6 82.5 23.6 46.9
MagR-LoRA $r=64$ 239M 38.6 60.5 81.1 24.8 51.2
LoftQ $r=64$ 239M 37.0 55.9 87.7 21.7 50.6
ApiQ-bw $r=64$ 239M 43.1 59.2 85.1 23.4 52.7
GPTQ-DiaBlo $N=64$ 189M 41.7 58.0 87.4 23.2 52.6
MagR-DiaBlo $N=64$ 189M 44.5 59.9 88.2 27.6 55.1
@@ 4.5 Memory and Computation Efficiency of DiaBlo
The simple structure of DiaBlo allows the use of efficient batched operations, thus reducing training time and memory usage. In this subsection, we evaluate the memory consumption and training speed of DiaBlo and compare them with other PEFT training methods.
@@
To further validate efficiency, we benchmark the fine-tuning speed of DiaBlo on the 170k commonsense reasoning dataset, comparing it against LoRA and DoRA. All experiments are conducted on a single A100 GPU with 80GB memory, using a batch size of 8 and a gradient accumulation step of 4. DiaBlo with $N=64$ and $N=128$ matches the training speed of LoRA with rank $r=32$, requiring 170 minutes per epoch, whereas DoRA is substantially slower at 480 minutes per epoch. These results confirm the practical efficiency of the proposed DiaBlo method.
4.6 Sufficiency of Diagonal Blocks
We investigate whether diagonal blocks provide a more effective subset for fine-tuning compared to alternative sparsity patterns. Following the methodology of Deng et al. (2024), we first analyze the performance retained from a fully fine-tuned model. Specifically, we fully fine-tune LLaMA3-8B on GSM8K and compute the update matrix $\Delta=\mathbf{W}_{\text{finetuned}}-\mathbf{W}_{\text{pretrained}}$. We then measure accuracy after reapplying only selected parts of $\Delta$, scaled by a factor $s$, to the pre-trained weights. We compare DiaBlo’s diagonal block selection with random entries, random blocks of the same size as DiaBlo, random columns, and random rows. As shown in Table 5, retaining diagonal blocks preserves accuracy substantially better than all other sparsity patterns, and degrades more gracefully as sparsity increases. We further evaluate these sparse patterns in direct fine-tuning. Table 5 shows that DiaBlo not only achieves the highest accuracy but also runs more efficiently than unstructured random methods, while avoiding the severe accuracy loss of row- or block-based updates. Together, these results demonstrate that diagonal blocks form a principled and hardware-friendly structure for sparse fine-tuning, sufficient to capture task-relevant updates.
Table 5: Comparison of DiaBlo with random sparse patterns on GSM8K. Left part reports accuracy from retaining sparse updates, and right part reports accuracy and training time from fine-tuning.
Method Retained Sparse Update Accuracy (%) Fine-tuning (Sparsity 1/64)
Sparsity 1/32 Sparsity 1/64 Sparsity 1/128 Accuracy (%) Time (min)
DiaBlo 67.30 65.47 63.11 67.68 17.26
Random Entries 65.98 64.05 62.19 65.35 26.51
Random Block 65.96 63.75 62.98 64.86 29.76
Random Column 65.73 62.93 61.66 65.19 17.01
Random Row 49.14 24.08 2.85 61.71 17.76